Racetrack A has 35,000 more grandstand seats than three times the number of grandstand seats at another race track B. Together, these two race tracks seat
287,000 auto racing fans. How many seats does each race track have?

The number of seats in race track A is _______
and the number of seats in race track B is _______

Answer by Cromlix(4381)

Hi there,
Make Racetrack B number of seats = 'x'
Racetrack A number of seats = 35,000 + 3x
x + (35,000 + 3x) = 287,000
x + 35,000 + 3x = 287,000
Collect like terms
x + 3x = 287,000 - 35,000
4x = 252,000
x = 252,000/4
x = 63,000.
Racetrack A has 224,000 seats
Racetrack B has 63,000 seats
Hope this helps :-)